# Reasoning Language Models: A Blueprint

### v1

@@ C.3.1 Datasets Generation and Preparation
Input: Policy Model $\pi_{\theta}$, tokenized dataset $D_{\text{SFT}}=\{(x^{(i)},y^{(i)})\}$, training hyperparameters (optimizer, learning rate $\eta$, batch size $B$, and maximum number of epochs $E$, …).
@@ C.3.2 SFT of the Policy
C.3.3 Q-Value Model Training
@@
By the end of training, $Q_{\varphi}$ should output accurate q-value scores, which will later guide policy refinement in Phase II and will improve the search accuracy when used in MCTS.
@@
Input: Q-Value Model $Q_{\varphi}$ (QVM), dataset $D_{\text{QVM-train}}=\{(x^{(i)},y^{(i)})\}$, training hyperparameters (optimizer, learning rate $\eta$, batch size $B$, and maximum epochs $E$, …).
@@
6:           Predict the q-value usind the QVM $\hat{y}^{(i)}=Q_{\varphi}(x^{(i)})$.
Compute the loss:
7:           if Regression Loss then
8:              $\mathcal{L}=\frac{1}{B}\sum_{(x^{(i)},y^{(i)})}(\hat{y}^{(i)}-y^{(i)})^{2}$.
9:           end if
10:           if Classification Loss then
11:              $\mathcal{L}=\frac{1}{B}\sum_{(x^{(i)},y^{(i)})}\texttt{BCE}(\hat{y}^{(i)},y^{(i)})$.
12:           end if
13:           Backpropagate gradients, update $\varphi$ via optimizer.
14:        end for
@@
C.4 Phase 2: RL Tuning of Policy with MCTS
@@
C.4.1 Phase 2 Algorithm
Notation. We present our notation with RL terminology which will be used in phase 2 training. For a reminder, the MCTS builds a tree of nodes that each define a reasoning sequence: for example $s_{i}=(z_{0},\dots,z_{i})$. These nodes represent a state. An action $a_{i}$ corresponds to the next reasoning step $z_{i+1}$ obtained when an LLM is queried with the state $s_{i}$ as context. Finally, a q-value estimate $q(s)$ is associated to each node $s$.
Thus, each reasoning sequence in the tree can be decomposed into a set of state, action and q-value $(s,a,q)$ triplets: $s^{(i)}_{t}=\{s^{(i)}_{j},z^{(i)}_{j+1},q(s^{(i)}_{j})\}_{j\leq t}$. These triplets can be used to perform reinforcement learning.
Phase 2 assumes an initial dataset $D_{q}=\{z_{0}^{(i)}\}$ of prompt questions - these questions may be isolated from the phase 1 dataset $D_{\text{SFT}}$. The training process (Algorithm 4) involves a repetition of a rollout phase followed by a training (reinforcement) phase.
In the rollout phase, an MCTS tree is constructed (Algorithm 1 in training mode) which generates sequences of reasoning steps. The individual states of the tree are attributed a value (see 1). These sequences are then stored in a replay buffer $\mathcal{R}$.
The reinforcement phase samples a batch of reasoning sequences from the replay buffer. From each trajectory, constituent states, actions and value estimatesand uses the corresponding values attributed during MCTS to perform RL training (for example with PPO or Reinforce). Alternative schemes may involve selecting preference pairs among trajectories and then aligning the policy using DPO, or simply selecting the most desirable trajectory per question and performing further SFT training.
During this reinforcement phase, the value model is also updated to mimic the (backpropagated) values from the MCTS process (Algorithm 7).
@@ C.4.1 Phase 2 Algorithm
Input: Policy $\pi_{\theta}$, q-value model $Q_{\varphi}$, dataset $D_{q}=\{z_{0}^{(i)}\}$, MCTS and RL hyperparameters.
Output: Trained $\pi_{\tilde{\theta}}$ and updated $Q_{\tilde{\varphi}}$.
@@
2:      ————– Rollout ———————————-
3:     for each question $z_{0}^{(i)}\in D_{q}$ do
4:        Generate MCTS tree with $\pi_{\theta}$ guided by $Q_{\varphi}$
(Algorithm 1)
5:        Extract individual trajectories and store them in a replay buffer =$\mathcal{R}=\{\tau^{(j)}\}$
6:     end for
7:      ————– Training ———————————
8:     for each epoch do
9:        Sample a batch $\mathcal{B}$ from replay buffer $\mathcal{R}$.
10:        Update policy $\pi_{\theta}$ (Algorithm 5).
11:        Update q-value model $Q_{\varphi}$ (Algorithm 7).
12:     end for
13:  end for
C.4.2 Policy Update
@@
$V(s_{t+1})=\frac{1}{\gamma}Q_{\varphi}(s_{t},a_{t})$
since rewards are sparse. The standard PPO approach trains the critic model from scratch on bootstrapped rewards for this purpose. We introduce an alternative advantage computation scheme that leverages the backpropagated values from Monte Carlo Tree Search (MCTS) in conjunction with $Q_{\psi}$, as detailed in Algorithm 6. This integration combines MCTS’s exploration and evaluation capabilities with the RL update, enhancing robustness and efficiency in reasoning tasks.
Further regularization can be imposed on the PPO training procedure. To align the policy $\pi_{\theta}$ with a reference policy $\pi_{\text{ref}}$ (usually instantiated as $\pi_{\theta}$ before phase 2) during training, the KL divergence $\texttt{KL}(\pi_{\theta}||\pi_{\text{ref}})$, between the two distributions can be added to the training loss. Additionally, to maintain the diversity of policy generations (and exploration during training), the entropy of the policy distribution can be enhanced by substracting it from the loss. We denote the additional entropy loss term over a batch $\mathcal{D}$ of state-action pairs $(s,a)$ as:
$\mathcal{L}_{H}=-\frac{1}{|\mathcal{D}|}\sum_{(s,a)\in\mathcal{D}}\sum_{a}\pi_{\theta}(a|s)\log\pi_{\theta}(a|s).$
Direct Preference Optimization (DPO). DPO aligns the policy to user preferences expressed as pairwise comparisons between reasoning sequences. Given pairs $(s^{+},s^{-})$, where $s^{+}$ is preferred over $s^{-}$. This method doesn’t require a reward/value model.
Supervised Fine-Tuning (SFT). As a straightforward alternative to RL, high-value reasoning sequences can be selected to perform SFT, i.e. train the policy to maximize the likelihood of these reasoning steps. This approach is inspired by AlphaZero-like frameworks, focusing on iteratively refining the policy to generate high-quality reasoning trajectories without requiring explicit rewards.
C.4.3 Advantage Calculation (for PPO Policy Updates)
@@
11:   ————– Train via DPO (pairwise preferences) ——-
@@
18:   ————– Train via SFT (single target sequence) ——-
@@
C.4.4 Q-Value Model Update
During phase 2, the q-value model $Q_{\varphi}$ is also updated to track MCTS-backtracked value estimates $q^{\text{MCTS}}(s_{t})$ which should be of higher quality (thanks to the final answer verifier and score aggregation from child nodes). For each subsequence $s_{t}=(z_{0},z_{1},\ldots,z_{t})$, we train the q-value model $Q_{\varphi}$ via squared error minimization, to match its values $Q_{\varphi}(s_{t})$ as closely as possible to the MCTS-values $Q^{\text{MCTS}}(s_{t})$.
@@ C.4.4 Q-Value Model Update
1:  for each $(s_{t},r_{s_{t}})\in\mathcal{D}$ do
2:     Compute loss:
$\mathcal{L}_{q}=\frac{1}{|\mathcal{D}|}\sum_{(s_{t},r_{s_{t}})}(Q_{\varphi}(s_{t})-q^{\text{MCTS}}(s_{t}))^{2}$.
3:  end for
4:  Perform gradient update on $\mathcal{L}_{q}$.
@@ D.2 Generating Data for Phase 1 Value Model Training
$V_{\pi_{\theta}}(s)=\mathbb{E}_{\tau\sim\pi_{\theta}}\left[\sum_{t=i}^{T}\gamma^{t-i}r(s_{t},a_{t})\mid s_{i}=s\right]$
@@
$V_{\pi_{\theta}}(s_{t})=\mathbb{E}_{\pi_{\theta}}\left[\gamma^{T-t}r(s_{T},a_{T})\mid s_{t}\right]$
@@
$V_{\pi_{\theta}}(s_{t})\approx\frac{1}{N}\sum_{i=1}^{N}\gamma^{T-t}r(s_{T}^{(i)},a_{T}^{(i)}):=\hat{V}(s_{t})$
@@ D.2.2 Continuous and Bounded Rewards: Modeling the Expected Future Reward
$V_{\pi_{\theta}}(s_{t})=\mathbb{E}\left[\gamma^{T-t}r(s_{T},a_{T})\mid s_{t}\right]$
@@
$V_{\pi_{\theta}}(s_{t})\approx\frac{1}{N}\sum_{i=1}^{N}\gamma^{T-t}r(s_{T}^{(i)},a_{T}^{(i)}):=\hat{V}(s_{t})$
@@
$\mathcal{L}(\varphi)=\frac{1}{N}\sum_{i=1}^{N}\left(\hat{V}_{\varphi}(s_{t}^{(i)})-\gamma^{T-t}r(s_{T}^{(i)},a_{T}^{(i)})\right)^{2}$
@@ D.3.2 The Difference between Value and Q-Value Models
$\hat{V}(s_{t+1})=\frac{1}{\gamma}\hat{Q}_{\varphi}(s_{t},a_{t})$

### v2

@@ C.2.1 Datasets Generation and Preparation
Input: Policy Model $\pi_{\theta}$, tokenized dataset $D_{\text{SFT}}=\{(x^{(i)},y^{(i)})\}$, training hyperparameters (optimizer, learning rate $\eta$, batch size $B$, and maximum number of epochs $E$.).
@@ C.2.2 SFT of the Policy
C.2.3 Q-Value Model Training
@@
By the end of training, $Q_{\varphi}$ should output accurate q-value scores, which will later guide policy refinement in Phase II and will improve the search accuracy when used in the MCTS.
@@
Input: Q-value model $Q_{\varphi}$ (QVM), dataset $D_{\text{QVM-train}}=\{(x^{(i)},y^{(i)})\}$, training hyperparameters (optimizer, learning rate $\eta$, batch size $B$, and maximum epochs $E$).
@@
6:           Predict the q-value with QVM $\hat{y}^{(i)}=Q_{\varphi}(x^{(i)})$.
7:           {Compute the loss:}
8:           if Regression Loss then
9:              $\mathcal{L}=\frac{1}{B}\sum_{(x^{(i)},y^{(i)})}(\hat{y}^{(i)}-y^{(i)})^{2}$.
10:           end if
11:           if Classification Loss then
12:              $\mathcal{L}=\frac{1}{B}\sum_{(x^{(i)},y^{(i)})}\texttt{BCE}(\hat{y}^{(i)},y^{(i)})$.
13:           end if
14:           Backpropagate gradients, update $\varphi$ via optimizer.
@@
17:  end for
C.3 Training Phase 2: RL Tuning of Policy with MCTS
@@
C.3.1 Phase 2 Algorithm
Phase 2 uses a set $D_{p}=\{p^{(i)}\}$ of prompt questions - these questions may be isolated from the phase 1 dataset $D_{\text{SFT}}$. The training process (Algorithm 4) involves a repetition of a MCTS rollout phase followed by a training (reinforcement) phase.
Data generation: MCTS rollout. To obtain data for the training, a MCTS tree $\mathcal{T}^{(i)}$ is build w.r.t. each question $p^{(i)}$ using the algorithm in Algorithm  1 in training mode. The set of hyperparameters for MCTS $\Xi_{MCTS}$, denotes the number of MCTS iterations $N$ (per question), the number of children expanded in every selection phase $M$, exploration constants $c_{1},c_{2}$, and backpropagation weight $\alpha$. To enhance the quality of the data, we prune the generated MCTS tree $\tilde{\mathcal{T}}^{(i)}=(\tilde{N}^{(i)},\tilde{E}^{(i)})$ to only include paths that reached a terminal state since only these paths received the reward. Then, we extract all nodes and a set of node characteristics from the pruned tree. The dataset comprises of state, action and q-value triplets of the pruned tree: $\{(s_{j}^{(i)},z_{j}^{(i)},q(s_{j}^{(i)})\}_{s_{j}\in\tilde{N}^{(i)}}$ The data is stored in a replay buffer $\mathcal{R}$.
The training: RL phase. The reinforcement phase samples a batch of reasoning sequences from the replay buffer. From each trajectory, constituent states, actions and value estimatesand uses the corresponding values attributed during MCTS to perform RL training (for example with PPO or Reinforce). Alternative schemes may involve selecting preference pairs among trajectories and then aligning the policy using DPO, or simply selecting the most desirable trajectory per question and performing further SFT training.
During this reinforcement phase, the value model is updated to mimic the (backpropagated) values from the MCTS process (Algorithm 7).
@@ C.3.1 Phase 2 Algorithm
Input: Policy $\pi_{\theta}$, q-value model $Q_{\varphi}$, dataset $D_{p}=\{p^{(i)}\}$, MCTS hyperparameters $\Xi_{MCTS}$.
Output: Trained $\pi_{\theta}$ and updated $Q_{\varphi}$.
@@
2:      ————– Rollout ———————————
3:     for each question $p^{(i)}\in D_{p}$ do
4:        {Generate MCTS tree with $\pi_{\theta}$ and $Q_{\varphi}$ (Algorithm 1)}
5:        $\mathcal{T}^{(i)}\leftarrow MCTS(p^{(i)},Q_{\varphi},\pi_{\theta},\Xi_{MCTS})$
6:        {Remove incomplete paths from the tree}
7:        $\tilde{\mathcal{T}}^{(i)}\leftarrow Prune(\mathcal{T}^{(i)})$
8:        {Extract nodes and values, store them in replay buffer}
9:        $\mathcal{R}\leftarrow\mathcal{R}\cup\{(s_{j}^{(i)},z_{j}^{(i)},q(s_{j}^{(i)})\}_{s_{j}\in\tilde{N}^{(i)}}$
10:     end for
11:      ————– Training ———————————
12:     for each epoch do
13:        Sample a batch $\mathcal{B}$ from replay buffer $\mathcal{R}$.
14:        Update policy $\pi_{\theta}$ (Algorithm 5).
15:        Update q-value model $Q_{\varphi}$ (Algorithm 7).
16:     end for
17:  end for
C.3.2 Policy Update
@@
$V(s_{t+1})=\frac{1}{\gamma}Q_{\varphi}(s_{t},a_{t}),$
since rewards are sparse. The standard PPO approach trains the critic model from scratch on bootstrapped rewards for this purpose. We introduce an alternative advantage computation scheme that leverages the backpropagated values from Monte Carlo Tree Search (MCTS) in conjunction with $Q_{\varphi}$, as detailed in Algorithm 6. This integration combines MCTS’s exploration and evaluation capabilities with the RL update, enhancing robustness and efficiency in reasoning tasks.
Further regularization can be imposed on the PPO training procedure. To align the policy $\pi_{\theta}$ with a reference policy $\pi_{\text{ref}}$ (usually instantiated as $\pi_{\theta}$ before phase 2) during training, the KL divergence $\texttt{KL}(\pi_{\theta}||\pi_{\text{ref}})$ , between the two distributions can be added to the training loss. Additionally, to maintain the diversity of policy generations (and exploration during training), the entropy of the policy distribution can be enhanced by subtracting it from the loss. The entropy penalty is estimated over a batch $\mathcal{D}$ of state-action pairs $(s,a)$ where $s$ denotes a reasoning sequence and $a$ the next reasoning step. The entropy of a single completion $a$ is computed by summing the entropy of its individual tokens $a_{1:|a|}$ of $a$:
$\mathcal{L}_{H}=-\frac{1}{|\mathcal{D}|}\sum_{(s,a)\in\mathcal{D}}\sum_{a_{i}\in a}\pi_{\theta}(a_{i}|[s,a_{1:{i-1}}])\log\pi_{\theta}(a|[s,a_{1:{i-1}}]).$
Direct Preference Optimization (DPO). DPO [115] aligns the policy to user preferences expressed as pairwise comparisons between reasoning sequences. Given pairs $(s^{+},s^{-})$, where $s^{+}$ is preferred over $s^{-}$. This method may not require a process reward/value model. The loss involves the sigmoid function which we denote as $\sigma$.
Supervised Fine-Tuning (SFT). As a straightforward alternative to RL, high-value reasoning sequences can be selected to perform SFT, i.e. train the policy to maximize the likelihood of these reasoning steps. The high-value reasoning sequences may be selected as terminal nodes having the highest q-value, or highest aggregated intermediate-step values. This approach is inspired by AlphaZero-like frameworks, focusing on iteratively refining the policy to generate high-quality reasoning trajectories without requiring explicit rewards.
C.3.3 Advantage Calculation (for PPO Policy Updates)
@@
11:   ————– Train via DPO (pairwise preferences) ——
@@
18:   ————– Train via SFT (single target sequence) ——
@@
C.3.4 Q-Value Model Update
During phase 2, the q-value model $Q_{\varphi}$ is also updated to track MCTS-backtracked value estimates $q^{\text{MCTS}}(s_{t})$ which should be of higher quality (thanks to the final answer verifier and score aggregation from child nodes). For each state-action pair $(s,a)$, we train the q-value model $Q_{\varphi}$ via squared error minimization, to match its q-value $Q_{\varphi}(s,a)$ as closely as possible to the corresponding MCTS-value $q^{\text{MCTS}}(s^{\prime})$ which saves the updated q-value of action $a$ taken in state $s$ leading to state $s^{\prime}$.
@@ C.3.4 Q-Value Model Update
1:  Compute loss:
$\mathcal{L}_{q}=\frac{1}{|\mathcal{D}|}\sum_{(s,a,s^{\prime})}(Q_{\varphi}(s,a)-q^{\text{MCTS}}(s^{\prime}))^{2}$.
2:  Perform gradient update on $\mathcal{L}_{q}$.
@@ D.2 Generating Data for Phase 1 Value Model Training
$V_{\pi_{\theta}}(s)=\mathbb{E}_{\tau\sim\pi_{\theta}}\left[\sum_{t=i}^{T}\gamma^{t-i}r(s_{t},a_{t})\mid s_{i}=s\right],$
@@
$V_{\pi_{\theta}}(s_{t})=\mathbb{E}_{\pi_{\theta}}\left[\gamma^{T-t}r(s_{T},a_{T})\mid s_{t}\right].$
@@
$V_{\pi_{\theta}}(s_{t})\approx\frac{1}{N}\sum_{i=1}^{N}\gamma^{T-t}r(s_{T}^{(i)},a_{T}^{(i)}):=\hat{V}(s_{t}),$
@@ D.2.2 Continuous and Bounded Rewards: Modeling the Expected Future Reward
$V_{\pi_{\theta}}(s_{t})=\mathbb{E}\left[\gamma^{T-t}r(s_{T},a_{T})\mid s_{t}\right],$
@@
$V_{\pi_{\theta}}(s_{t})\approx\frac{1}{N}\sum_{i=1}^{N}\gamma^{T-t}r(s_{T}^{(i)},a_{T}^{(i)}):=\hat{V}(s_{t}),$
@@
$\mathcal{L}(\varphi)=\frac{1}{N}\sum_{i=1}^{N}\left(\hat{V}_{\varphi}(s_{t}^{(i)})-\gamma^{T-t}r(s_{T}^{(i)},a_{T}^{(i)})\right)^{2}.$
@@ D.3.2 The Difference between Value and Q-Value Models
$\hat{V}(s_{t+1})=\frac{1}{\gamma}\hat{Q}_{\varphi}(s_{t},a_{t}).$